# Benchmarking Deep Neural Networks for Modern Recommendation Systems

### v1

@@ IV Experimentation and Results
To guide our experimental protocol and evaluate the effectiveness of the seven neural network architectures, the following research questions (RQs) have been formulated:
  • RQ1: How do different neural network architectures perform in terms of accuracy across heterogeneous recommendation datasets?
  • RQ2: Which models provide the highest recall and F1-score, and how consistent are these performances across datasets?
  • RQ3: How does the top-$k$ accuracy degrade as $k$ increases, and which models retain the highest performance as the recommendation list expands?
  • RQ4: Which neural network architecture provides the greatest intra-list diversity in top-$k$ recommendations?
  • RQ5: How do accuracy and diversity interact within each model, and which architectures offer the best trade-off between relevance and variety?
These research questions are addressed systematically throughout the following subsections.
IV-A Dataset (RQ1)
In our research, we have chosen three distinct datasets from different e-commerce and streaming platforms to evaluate the strength and broad applicability of our methods. These datasets were preprocessed as outlined in earlier sections.
Retail Rocket Dataset: This dataset includes category trees, item properties, and user behavior logs, providing insights into consumer actions such as visits and checkouts.
Amazon Product Dataset: Sourced from Amazon, it contains product metadata and user reviews, offering rich information on item popularity and user-item interactions.
Netflix Prize Dataset: This dataset includes detailed user ratings for movies, making it suitable for evaluating temporal user preferences and recommendation quality in streaming platforms.
IV-B Evaluation Metrics (RQ1, RQ2)
To assess the performance of the neural networks, we adopt widely recognized evaluation metrics:
Accuracy:
@@ IV-B Evaluation Metrics (RQ1, RQ2)
IV-C Justification of Model Parameters (RQ1)
The neural network configurations were selected after iterative tuning to achieve optimal performance.
@@ IV-C Justification of Model Parameters (RQ1)
IV-D Overview of Model Performance (RQ1, RQ2)
Performance across the three datasets is summarized in Tables 3–5 and visualized in Figures 2–4.
TABLE III: Performance metrics on Retail Rocket E-commerce dataset
Model Accuracy (%) Recall (%) F1-Score (%)
CNN 80 83 80.5
RNN 83 86 83.5
GNN 88 91 88.5
Autoencoder 73 76 73.5
Transformer 86 88 86.5
NCF 82 81 82.5
Siamese Networks 81 83 81.5
TABLE IV: Performance metrics on Amazon dataset
Model Accuracy (%) Recall (%) F1-Score (%)
CNN 78 79 77.5
RNN 82 85 82.5
GNN 90 93 90.5
Autoencoder 70 73 70.5
Transformer 88 90 88
NCF 85 84 85.5
Siamese Networks 84 86 84.5
TABLE V: Performance metrics on Netflix Prize dataset
Model Accuracy (%) Recall (%) F1-Score (%)
CNN 75 78 75.5
RNN 89 92 89.5
GNN 85 87 85.5
Autoencoder 72 75 72.5
Transformer 83 85 83.5
NCF 80 79 80
Siamese Networks 78 80 78.5
Figure 2: Metrics visualization on Retail Rocket E-commerce dataset
Figure 3: Metrics visualization on Amazon dataset
Figure 4: Metrics visualization on Netflix Prize
IV-E Accuracy and Diversity Evaluation on Top-$k$ (RQ3, RQ4)
Tables 6–12 present accuracy and intra-list diversity for top-$k$ recommendations for each model-dataset pair. These results are visualized in Figure 5.
TABLE VI: CNN performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 86.12 0 84.76 0 82.66 0
@@ IV-E Accuracy and Diversity Evaluation on Top-$k$ (RQ3, RQ4)
TABLE VII: RNN performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 89.43 0 88.48 0 93.41 0
@@
TABLE VIII: GNN performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 95.63 0 98.47 0 88.39 0
@@
TABLE IX: Autoencoder performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 80.15 0 76.14 0 76.14 0
@@
TABLE X: Transformer performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 90.43 0 97.28 0 85.39 0
@@
TABLE XI: NCF performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 89.31 0 96.18 0 83.23 0
@@
TABLE XII: Siamese Networks performance of top-$k$ recommendations on all datasets
Top $k$ Retail E-commerce Amazon Products Netflix Prize
Accuracy (%) ILD (%) Accuracy (%) ILD (%) Accuracy (%) ILD (%)
1 87.21 0 94.32 0 84.26 0
@@
Figure 5: Accuracy and intra-list diversity visualization on top k recommendations
IV-F Interpretation of Results (RQ5)
The evaluation of the seven architectures—CNN, RNN, GNN, Autoencoder, Transformer, NCF, and Siamese Networks—reveals substantial performance variations based on dataset characteristics.
GNNs achieve the highest accuracy in the Retail E-commerce and Amazon datasets due to their ability to model complex graph structures. Transformers also perform strongly due to their capacity to extract global relationships through attention.
For the Netflix dataset, RNNs outperform others by effectively capturing sequential patterns in user rating histories, while GNNs remain robust but slightly less suited to temporal dynamics.
In terms of diversity, Siamese Networks offer the highest ILD on Retail Rocket due to their similarity-based optimization. CNNs lead in Amazon due to feature extraction variety, while GNNs yield the most diverse results for Netflix due to rich graph connectivity.
These results indicate that:
  • GNNs and Transformers are optimal for accuracy-focused recommendations.
  • Siamese Networks and CNNs provide strong diversity and can enhance hybrid recommenders.
  • The best model depends on the desired trade-off between accuracy and diversity.
IV-G Comparison with Other Works (RQ1, RQ2, RQ4)
Research on recommendation systems often relies on deep learning models such as matrix factorization techniques and conventional neural networks. Our approach, leveraging advanced models such as Graph Neural Networks (GNNs) and Transformers, brings a cutting-edge perspective by incorporating recent neural network developments. For instance, GNNs effectively model complex inter-item relationships, while Transformers utilize attention mechanisms, significantly enhancing traditional approaches.
While many conventional methods primarily focus on accuracy, our strategy emphasizes the diversity of recommendations. The Siamese Network, in particular, provides a unique contribution by emphasizing item-to-item similarities, broadening user engagement through more varied recommendations. This contrasts with typical systems that may overlook the importance of diversity on user satisfaction.
Our models also incorporate temporal dynamics, often underexplored in prior research. Recurrent Neural Networks (RNNs) excel in sequence prediction, allowing our systems to adapt recommendations to evolving user behavior. This adaptability is especially crucial for platforms like Netflix, where user preferences can change rapidly.
Across key metrics such as accuracy, recall, and F1-score, our models consistently demonstrate strong performance. The GNN model achieves superior results on the Retail Rocket and Amazon datasets, while the RNN model excels on the Netflix dataset. These outcomes surpass many existing solutions, providing more robust and reliable recommendations.
Furthermore, the effectiveness of our models in handling complex datasets, such as Amazon Products and Netflix Prize, demonstrates their robustness. These datasets contain a broad range of items and highly subjective user preferences, posing substantial challenges that our models successfully navigate.
Our comparative analysis highlights the potential for hybrid approaches that combine the strengths of various architectures to optimize both accuracy and diversity. Moreover, the results advocate for continued exploration of graph-based and attention-driven models in complex recommendation scenarios, with applications extending beyond e-commerce and media content.
Our models not only meet but advance current standards in recommendation system research. They provide a solid foundation for future efforts aimed at enhancing the scalability and adaptability of recommendation engines to diverse and dynamic market conditions.
IV-H Discussion on Limitations and Complications (RQ5)
Despite the advances achieved by our neural network models, several limitations and complications have been identified, highlighting areas for future improvement.
Computational Complexity: Advanced models like GNNs and Transformers require substantial computational resources due to their complex architectures. This can result in extended training times and higher operational costs, which may be impractical in environments with limited hardware or where rapid response is critical.
Dependency on Large, High-Quality Datasets: The models’ performance depends heavily on the availability of extensive and clean data. Sparse, noisy, or incomplete datasets can lead to reduced accuracy and biased recommendations. Additionally, reliance on large datasets raises concerns about user privacy and data security.
Balancing Accuracy and Diversity: Although our models perform well in complex environments, achieving a balance between recommendation accuracy and diversity remains challenging. Highly accurate recommendations may not always be sufficiently diverse, potentially limiting user engagement. Fine-tuning the models to recommend a broader range of items without compromising personalization is an ongoing challenge.
Adaptability to Real-World Dynamics: While our models incorporate temporal changes in user preferences, more dynamic real-time adaptation is needed to respond to sudden shifts in trends or user behavior. This is particularly relevant in fast-evolving domains such as fashion or entertainment, where user interests can change quickly.
While our models represent a significant advancement in recommendation systems, these limitations underscore the need for ongoing refinement. Future work should aim to improve computational efficiency, enhance data handling capabilities, and optimize the balance between diversity and relevance in real-world dynamic environments.
@@ V Conclusion
Throughout this study, we have rigorously evaluated the performance of seven advanced neural network models across three diverse datasets—Retail E-commerce, Amazon Products, and Netflix Prize. Our findings demonstrate that these models, particularly Graph Neural Networks (GNNs) and Transformers, provide significant improvements over traditional recommendation systems by effectively handling complex data structures and emphasizing both accuracy and diversity in recommendations.
The GNN model excels in environments with intricate inter-item relationships, capturing the rich connectivity within user-item interactions. Transformers, with their efficient attention mechanisms, handle large datasets effectively, discerning nuanced relationships within sequences. Meanwhile, RNNs show superior performance in capturing temporal dynamics, making them especially suitable for platforms like Netflix, where user preferences evolve rapidly. This variation in model efficacy highlights the potential of tailored approaches that align with the specific characteristics and requirements of each dataset.
Despite these promising results, the study identifies several limitations, including computational demands, dependency on large, high-quality datasets, and challenges in balancing recommendation accuracy with diversity. These limitations underscore the need for ongoing research to optimize neural network architectures, improve computational efficiency, and enhance data processing strategies.
Looking ahead, the integration of hybrid models that combine the strengths of multiple architectures could lead to more robust, adaptable, and user-centric recommendation systems. Additionally, addressing ethical considerations related to data privacy and developing mechanisms capable of dynamically adjusting to real-world changes will be critical for the continued relevance of these systems.
This research not only advances our understanding of neural network applications in recommendation systems but also lays the groundwork for future innovations that can transform the way users interact with digital platforms. By continually refining these models, we can better meet the ever-changing preferences of users, ensuring that recommendation systems remain both relevant and impactful in the digital age.
Author Contributions
Abderaouf Bahi: Writing – original draft, Validation, Methodology, Investigation, Conceptualization, Visualization, Software. Ibtissem Gasmi: Writing – review and editing, Methodology, Supervision.
Acknowledgements
The authors acknowledge the Algerian Ministry of Higher Education and Scientific Research (MESRS).

### v2

@@ IV Experimentation and Results
To evaluate neural architectures under the ROB framework, the following research questions are defined:
  • RQ1 (R1 – Predictive Accuracy): How do different neural network architectures perform in terms of accuracy across heterogeneous recommendation datasets?
  • RQ2 (R1 – Accuracy Robustness): Which models provide the highest recall and F1-score, and how consistent are these performances across datasets?
  • RQ3 (R1/R4 – Ranking Robustness): How does top-$k$ recommendation performance degrade as the recommendation list expands, and which architectures maintain stable performance at larger $k$ values?
  • RQ4 (R2 – Recommendation Diversity): Which neural network architecture provides the greatest intra-list diversity in top-$k$ recommendations?
  • RQ5 (R1–R2 Trade-off): How do predictive accuracy and recommendation diversity interact within each model, and which architectures achieve the most favorable balance between relevance and diversity?
Note: Requirements R3 (Relational Awareness) and R5 (Computational Efficiency) are evaluated through comparative performance analysis on graph-structured datasets and explicit measurements of training time, inference latency, and memory consumption, respectively, as reported in the subsequent experimental and results sections.
These research questions are addressed systematically in the following subsections, with results interpreted in relation to the core system requirements defined by the ROB framework.
IV-A Dataset and Applications Contexts
Three real-world datasets from distinct application domains are used to evaluate the generality and robustness of the proposed benchmark. All datasets are preprocessed following the procedure described in Section III to ensure consistency across experiments.
Retail Rocket Dataset: This dataset contains category hierarchies, item attributes, and user interaction logs, capturing consumer behavior such as product views and checkouts. It reflects typical e-commerce interaction patterns.
Amazon Product Dataset: Sourced from Amazon, this dataset includes product metadata and user-generated reviews, providing rich signals of item popularity and user–item interactions.
Netflix Prize Dataset: This dataset consists of timestamped user ratings for movies and is well suited for analyzing temporal user preferences and recommendation performance in streaming platforms.
IV-B Evaluation Metrics
To quantitatively assess model performance under the ROB framework, we employ widely adopted evaluation metrics that capture predictive accuracy and recommendation diversity. These metrics are formally defined below and are used consistently across all experiments. Accuracy:
@@ IV-B Evaluation Metrics
IV-C Justification of Model Parameters
Model configurations were selected through iterative tuning to ensure stable convergence and competitive performance across datasets, while maintaining consistency with the unified evaluation protocol. The following summarizes the architectural choices and key parameters for each model.
@@ IV-C Justification of Model Parameters
IV-D Predictive Accuracy and Robustness Across Datasets
Tables IVVVI, visualized in Figures 34, summarize predictive accuracy, recall, and F1-score for all evaluated architectures across the Retail Rocket, Amazon, and Netflix Prize datasets. Under the ROB framework, these results correspond to Requirement R1 (Predictive Accuracy) and provide insight into the robustness of each model across heterogeneous application contexts. GNNs consistently achieve the highest performance on the Retail Rocket and Amazon datasets, reflecting their ability to leverage relational item– item structures common in e-commerce environments. In contrast, RNN exhibit the strongest performance on the Netflix Prize dataset, highlighting their effectiveness in modeling sequential user behavior in temporally driven recommendation scenarios. Transformer-based models also demonstrate competitive accuracy across datasets, particularly in leading positions where long-range dependencies are prominent. Autoencoder-based approaches show comparatively lower performance across all datasets, suggesting limitations in capturing complex interaction patterns when used in isolation. Overall, these findings indicate that no single architecture universally dominates across datasets, reinforcing the importance of selecting models in accordance with specific system requirements.
TABLE IV: Performance metrics on Retail Rocket E-commerce dataset
Model Accuracy(%) Recall(%) F1-Score(%) TT(s)
CNN 80 83 80.5 420
RNN 83 86 83.5 680
GNN 88 91 88.5 910
Autoencoder 73 76 73.5 360
Transformer 86 88 86.5 1150
NCF 82 81 82.5 290
Siamese 81 83 81.5 520
TABLE V: Performance metrics on Amazon dataset
Model Accuracy(%) Recall(%) F1-Score(%) TT(s)
CNN 78 79 77.5 2 850
RNN 82 85 82.5 4 200
GNN 90 93 90.5 5 800
Autoencoder 70 73 70.5 2 100
Transformer 88 90 88 7 900
NCF 85 84 85.5 1 750
Siamese 84 86 84.5 3 100
TABLE VI: Performance metrics on Netflix Prize dataset
Model Accuracy(%) Recall(%) F1-Score(%) TT(s)
CNN 75 78 75.5 1 120
RNN 89 89 89.5 1 950
GNN 85 87 85.5 2 600
Autoencoder 72 75 72.5 980
Transformer 83 85 83.5 3 400
NCF 80 79 80 820
Siamese 78 80 78.5 1 480
TABLE VII: Performance of top-$k$ recommendations across all datasets for all models
(a) (a) CNN
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 86.12 0.00 84.76 0.00 82.66 0.00
@@ IV-D Predictive Accuracy and Robustness Across Datasets
(b) (b) RNN
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 89.43 0.00 88.48 0.00 93.41 0.00
@@
(c) (c) GNN
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 95.63 0.00 98.47 0.00 88.39 0.00
@@
(d) (d) Autoencoder
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 80.15 0.00 76.14 0.00 76.14 0.00
@@
(e) (e) Transformer
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 90.43 0.00 97.28 0.00 85.39 0.00
@@
(f) (f) NCF
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 89.31 0.00 96.18 0.00 83.23 0.00
@@
(g) (g) Siamese Networks
Top-$k$ Retail Amazon Netflix
Acc ILD Acc ILD Acc ILD
1 87.21 0.00 94.32 0.00 84.26 0.00
@@
IV-E Top-k Performance Degradation and Diversity Trade-offs
Table VII and Figure 4 illustrate the evolution of predictive accuracy and intra-list diversity as the recommendation list size increases from top-1 to top-10. Under the ROB framework, these results jointly reflect Requirement R1 (predictive accuracy), Requirement R2 (recommendation diversity), and Requirement R4 (ranking robustness). Across all architectures and datasets, accuracy generally decreases as k increases, indicating the expected trade-off between recommendation breadth and relevance. In contrast, intra-list diversity consistently increases with larger k values, reflecting broader item exposure. GNN and RNN exhibit comparatively slower accuracy degradation while maintaining moderate diversity growth, suggesting greater robustness in expanding recommendation scenarios. Siamese Networks demonstrate the strongest diversity gains across datasets, particularly at larger k values, albeit with a more pronounced decline in accuracy. Autoencoder-based approaches show higher initial accuracy in small k settings but experience sharper degradation as k increases, indicating limited robustness in extended recommendation lists. These observations highlight that different architectures prioritize distinct objectives, reinforcing the multi-objective nature of modern recommendation systems.
Figure 2: Performance comparison across datasets using multiple evaluation metrics.
Figure 3: Performance comparison across datasets using multiple evaluation metrics.
(a) Retail Rocket E-commerce dataset
(b) Amazon dataset
(c) Netflix Prize dataset
Figure 4: Performance metrics visualization across three datasets: (a) Retail Rocket E-commerce, (b) Amazon, and (c) Netflix Prize.
IV-F Computational Efficiency Analysis
TABLE VIII: Computational Cost Analysis of Recommendation Models Across Datasets
Model Retail Rocket Amazon Products Netflix Prize
Train (s) Infer (ms) Mem (MB) Train (s) Infer (ms) Mem (MB) Train (s) Infer (ms) Mem (MB)
CNN 420 6.2 380 2 850 9.5 820 1 120 7.1 510
RNN 680 9.8 420 4 200 14.3 960 1 950 11.6 620
GNN 910 12.4 540 5 800 18.7 1 200 2 600 15.2 780
Autoencoder 360 5.1 310 2 100 7.8 650 980 6.3 430
Transformer 1 150 21.6 880 7 900 35.4 1 650 3 400 27.9 1 120
NCF 290 4.6 270 1 750 6.9 580 820 5.4 390
Siamese Networks 520 8.2 360 3 100 12.1 740 1 480 9.7 560
Table VIII reports the training time, inference latency, and memory consumption of all evaluated architectures across the three datasets, addressing Requirement R5 (Computational Efficiency) of the ROB framework. The results reveal substantial variation in computational cost across model classes, reflecting differences in architectural complexity and scalability.
Lightweight architectures such as NCF and Autoencoders exhibit the lowest training times, inference latency, and memory usage across all datasets, indicating favorable efficiency characteristics for large-scale or resource-constrained deployments. CNN-based models demonstrate moderate computational cost, offering a balance between efficiency and representational capacity. In contrast, GNNs and Transformer-based models incur significantly higher computational overhead, particularly on the Amazon Products and Netflix Prize datasets. This increase is attributed to graph message passing operations in GNNs and the quadratic complexity of self-attention mechanisms in Transformers. Recurrent Neural Networks and Siamese Networks occupy an intermediate position, with higher costs than convolutional and matrix-factorization-based approaches but lower overhead than attention- or graph-based architectures.
Overall, these findings highlight a clear trade-off between model expressiveness and computational efficiency. While complex architectures deliver strong predictive performance and richer relational modeling, their deployment may be constrained by training cost, inference latency, or memory requirements. This analysis underscores the importance of incorporating efficiency considerations alongside accuracy and diversity when selecting or combining models for real-world recommendation systems.
IV-G Interpretation of Results
The comparative evaluation of the seven neural architectures reveals that model performance is strongly influenced by dataset characteristics and underlying interaction patterns. Interpreted through the ROB framework, the results indicate that different architectures exhibit complementary strengths rather than universal dominance.
GNN achieve superior predictive accuracy on the Retail Rocket and Amazon datasets, reflecting their ability to exploit complex item–item relationships (R3) common in e- commerce environments. Transformer-based models also demonstrate competitive accuracy due to their capacity to capture global interaction patterns through attention mechanisms. In contrast, RNN outperform other architectures on the Netflix Prize dataset, highlighting their effectiveness in modeling sequential user behavior and temporal preference evolution.
With respect to recommendation diversity, Siamese Networks consistently yield higher intra-list diversity, particularly in retail-oriented datasets, owing to their similarity-based learning objective. CNN-based models also exhibit strong diversity characteristics in the Amazon dataset, while GNNs provide balanced diversity on Netflix due to rich relational connectivity. Overall, these findings emphasize that the choice of architecture should be guided by the desired balance between predictive accuracy and recommendation diversity, reinforcing the multi-objective nature of modern recommendation systems. Based on the observed performance trends across datasets and evaluation requirements, Table IX summarizes the empirical association between model classes and system requirements. It
TABLE IX: Empirical alignment between neural network architectures and system requirements under the ROB framework. Avg $T$: Average training time.
Model R1 R2 R3 R4 R5 (Avg. TT)
CNN Moderate High Low Moderate Low
RNN High Low High Low Moderate
GNN High Moderate High High High
Autoencoder Moderate Low Moderate Low Low
Transformer High Moderate Low High High
NCF High Low Low Low Low
Siamese Network Moderate Low High High Low
highlights that different neural architectures satisfy distinct system requirements under the ROB framework, with no single model excelling across all dimensions. This observation suggests that robust recommendation systems can benefit from ensemble or hybrid designs that integrate complementary modeling capabilities. For example, accuracy-oriented architectures such as GNNs, Transformers, or NCF can be combined with diversity-promoting models such as Siamese Networks or CNNs, while sequence-aware models like RNNs can enhance temporal adaptability. Such multi-model strategies provide a principled way to jointly address predictive accuracy, diversity, relational awareness, and temporal dynamics, offering a promising direction for building resilient and adaptable recommendation systems aligned with evolving application demands.
IV-H Comparison with Related Work
Existing research on recommendation systems has extensively explored deep learning approaches, including matrix factorization, conventional neural networks, and, more recently, graph-based and attention-based models. Prior studies typically evaluate these architectures in isolation or focus primarily on predictive accuracy. In contrast, the present study adopts a requirement-oriented evaluation perspective that jointly considers accuracy, diversity, relational modeling capability, and temporal robustness. While GNN and Transformers have previously been shown to perform well in recommendation tasks, our findings further contextualize their strengths under a unified evaluation protocol across heterogeneous datasets. Similarly, the diversity-promoting behavior of Siamese Networks, which has been less emphasized in earlier benchmarking studies, is highlighted through explicit diversity analysis.
By systematically evaluating multiple architectures under consistent experimental conditions, this work complements existing literature by providing comparative insights that inform architecture selection and combination, rather than proposing alternative models or optimization strategies.
IV-I Limitations
Despite the promising results obtained under the ROB framework, several limitations should be acknowledged. Although this study reports quantitative measurements of training time, inference latency, and memory consumption (R5), the computational efficiency analysis is conducted under a fixed hardware and implementation setting. As a result, absolute efficiency values may vary across deployment environments, software optimizations, or hardware accelerators. Nevertheless, the reported results provide a consistent relative comparison of computational cost across architectures. Model performance is also strongly influenced by data availability and quality. Sparse, noisy, or incomplete interaction data can degrade predictive accuracy and amplify bias, while large-scale data collection raises concerns related to privacy preservation and data security. Although sequence-aware architectures address temporal dynamics to some extent, the evaluated models are trained in offline settings and do not fully support real-time adaptation to abrupt shifts in user behavior or rapidly evolving trends, which remains a challenge in highly dynamic application domains.
In addition, achieving an effective balance between predictive accuracy (R1) and recommendation diversity (R2) remains non-trivial. Accuracy-oriented architectures may produce overly homogeneous recommendations, whereas diversity-focused approaches can reduce relevance. Addressing this trade-off requires further investigation into adaptive and hybrid strategies capable of dynamically adjusting system objectives based on contextual or user-driven signals. These limitations point toward future research directions that emphasize deployment-aware efficiency evaluation, robust data handling, and multi-objective optimization in real-world recommendation systems.
@@ V Conclusion
This paper presented a requirement-oriented benchmarking study of seven neural network architectures across three heterogeneous datasets, Retail E-commerce, Amazon Products, and Netflix Prize. By organizing evaluation around core system requirements, including predictive accuracy, recommendation diversity, relational awareness, temporal dynamics, and computational efficiency, the proposed ROB framework offers a unified and application-relevant perspective on modern recommendation system evaluation. Experimental results show that no single architecture consistently satisfies all requirements. GNNs and Transformers achieve strong predictive accuracy in settings characterized by complex relational structures, while RNNs are particularly effective for modeling temporal user behavior. In contrast, Siamese Networks and CNN-based models contribute more prominently to recommendation diversity, emphasizing broader item exposure. These findings highlight the importance of aligning model selection with specific system objectives and data characteristics. The analysis further reveals inherent trade-offs between model expressiveness and computational efficiency, as well as between predictive accuracy and recommendation diversity. Such trade-offs underscore the limitations of relying on a single architecture in real-world deployments and motivate the consideration of hybrid and ensemble strategies that integrate complementary modeling capabilities. Overall, this work contributes a structured benchmarking perspective that complements existing recommendation system research by emphasizing multi-objective evaluation and practical design implications. The ROB framework provides a reusable reference for evaluating and designing recommendation systems that better align with the evolving demands of contemporary digital platforms. Looking ahead, we will explore adaptive ensemble mechanisms that dynamically balance accuracy, diversity, and efficiency under changing user and system constraints. Future work may extend the ROB framework to evaluate large language model–based recommendation systems, where conversational reasoning, contextual understanding, and efficiency constraints introduce new evaluation challenges beyond traditional neural architectures.
Acknowledgment
The authors would like to thank the editors and anonymous reviewers for their time and valuable comments.